if the area of a rectangle with width x can be represented with the expression A(x)=x(14-x). what is the perimeter of the rectangle?
A.) 28
B.)56
C.) 56-4x
D.) 4x+28

Answer by Mathtut(3670)

A=L*W (x)(14-x)
:
P=2L+2W
:
P=2x+2(14-x)=2x+28-2x=28
:
answer is A